Tim and Judy mix two kinds of feed for pedigree dogs. They wish to make 27lbs of feed worth $0.46 per lb by mixing one kind worth $0.31 per lb with another worth $0.77 per lb. How many lbs of the cheaper kind should they use in the mix?
Need assistance on setting this problem up.

Answer by mananth(16946)

Pedigree food
Price ---------------- quantity
Type I 0.31 ---------------- x pounds
Type II 0.77 ------ 27 - x pounds
Mixture 0.46 ---------------- 27

0.31 x + 0.77 ( 27 - x ) = 12.42
0.31 x + 20.79 - 0.77 x = 12.42
0.31 x - 0.77 x = 12.42 - -20.79
-0.46 x = -8.37
/ -0.46
x = 18.2 pounds 0.31 Type I
8.8 pounds 0.77 Type II

m.ananth@hotmail.ca